### v2

# How Can We Effectively Expand the Vocabulary of LLMs with 0.01GB of Target Language Text?

Abstract
Large language models (LLMs) have shown remarkable capabilities in many languages beyond English. Yet, LLMs require more inference steps when generating non-English text due to their reliance on English-centric tokenizers and vocabulary, resulting in higher usage costs to non-English speakers. Vocabulary expansion with target language tokens is a widely used cross-lingual vocabulary adaptation approach to remedy this issue. Despite its effectiveness in inference speedup, previous work on vocabulary expansion has focused on high-resource settings assuming access to a substantial amount of target language data to effectively initialize the embeddings of the new tokens and adapt the LLM to the target language. However, vocabulary expansion in low-resource settings has yet to be explored. In this paper, we investigate vocabulary expansion in low-resource settings by considering embedding initialization methods and continual pre-training strategies. Through extensive experiments across typologically diverse languages, tasks and models, we establish a set of strategies to perform vocabulary expansion for faster inference, maintaining competitive downstream performance to baselines with only 30K sentences ($\sim$0.01GB text data) from the target language. [^1] Our code and models are available on GitHub.
How Can We Effectively Expand the Vocabulary of LLMs
with 0.01GB of Target Language Text?
Atsuki Yamaguchi1, Aline Villavicencio1,2  and Nikolaos Aletras1 1School of Computer Science, University of Sheffield, United Kingdom 2Department of Computer Science, Institute of Data Science and Artificial Intelligence, University of Exeter, United Kingdom {ayamaguchi1,a.villavicencio,n.aletras}@sheffield.ac.uk
Figure 1: We explore various adaptation strategies for cross-lingual vocabulary expansion (Refer to caption) that allow inference speedups while retaining competitive performance to the source models in low-resource settings.
1 Introduction
Large language models (LLMs) have strong capabilities in English and other languages OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Groeneveld et al. (2024). Yet, processing non-English texts with LLMs is challenging. They suffer from tokenization overfragmentation [^2] See Appendix E.1 for a quantitative analysis. and thus require more inference steps due to the reliance on English-centric tokenizers and vocabulary, resulting in higher utility costs for non-English speakers Ahia et al. (2023); Petrov et al. (2023).
Cross-lingual vocabulary adaptation (CVA) via vocabulary expansion has been proposed to adapt LLMs (including their tokenizers) to specific target languages. Vocabulary expansion approaches extend the vocabulary of a source model with tokens from a target language, followed by continual pre-training on target language data. A wide range of language-specific LLMs derived from an English-centric LLM such as Llama2 Touvron et al. (2023) have been made available following this approach, including Chinese Cui et al. (2023), Tamil Balachandran (2023), Portuguese Larcher et al. (2023), and Japanese Fujii et al. (2024) models, inter alia. Vocabulary expansion improves inference speed but often assumes access to a substantial amount of target language data for adaptation. For example, Chinese and Tamil LLaMAs make use of 20 and 12GB of target language text, respectively Cui et al. (2023); Balachandran (2023). Given the guaranteed large number of model updates in high-resource settings, the embeddings of new tokens are often simply randomly initialized, followed by the standard continual pre-training with a causal language modeling (clm) objective Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024). However, it is not clear how effective this approach to vocabulary expansion is in low-resource settings.
In this paper, we seek to (i) answer if this widely used adaptation approach under high-resource settings is as effective in low-resource settings; and (ii) identify the best possible vocabulary expansion strategies for language adaptation in low-resource settings, maintaining similar performance to the source models with faster inference (Figure 1). Our key contributions are as follows:
  • We present the first systematic study of vocabulary expansion-based adaptation of generative LLMs (i.e three English-centric models) in low-resource settings (i.e. assuming only 30K sentences, $\sim$0.01GB text data), on two generation tasks (i.e. machine translation and summarization) across ten typologically diverse languages.
  • Our results show that the popular vocabulary expansion approach in high-resource settings (i.e. random initialization and fine-tuning the full model with a clm objective) is not always optimal in low-resource settings.
  • We find that vocabulary initialization approaches that use heuristics information from the source and target tokinizers are more effective. Furthermore, fine-tuning the top and bottom two layers of the LLM using a multi-token prediction objective Gloeckle et al. (2024) works better than fine-tuning the full model with clm. Finally, using a short input sequence length by splitting longer text into multiple sentences allows for a larger number of model updates, mitigating underfitting.
2 Related Work
Text Overfragmentation.
LLMs tend to overfragment text in underrepresented languages Rust et al. (2021); Muller et al. (2021). Overfragmentation has significant implications for non-English speakers, including higher API costs Ahia et al. (2023); Petrov et al. (2023), slower inference Hofmann et al. (2022); Sun et al. (2023); Petrov et al. (2023), and lower downstream performance Bostrom and Durrett (2020); Rust et al. (2021); Toraman et al. (2023); Fujii et al. (2023).
Cross-lingual Vocabulary Adaptation.
CVA methods have been proposed for adapting base LLMs to specific target languages for improving downstream performance and inference speed Cui et al. (2023); Balachandran (2023); Fujii et al. (2024); Yamaguchi et al. (2024). CVA via vocabulary expansion incorporates new tokens into the source vocabulary Chau et al. (2020); Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Fujii et al. (2024); Liu et al. (2024). CVA with vocabulary replacement replaces the entire source vocabulary with a new one from the target Minixhofer et al. (2022); Dobler and de Melo (2023); Ostendorff and Rehm (2023); Downey et al. (2023). More recent methods extend only the language modeling head Hong et al. (2024). CVA is followed by continual pre-training on target language data, often called language adaptive pre-training – LAPT Chau et al. (2020).
Vocabulary expansion is widely used for developing language-specific generative LLMs from a source model, e.g. Chinese Cui et al. (2023) and Tamil Balachandran (2023) LLaMAs, and Swallow for Japanese Fujii et al. (2024). This line of work assumes access to a substantial amount of target language data, e.g. 312B characters of Japanese text used for Swallow. This might not be feasible in low-resource settings with limited target language data or compute. To the best of our knowledge, our work is the first to investigate vocabulary expansion for efficient LLM inference in extremely low-resource settings by assuming access to a small amount of target language data.
3 Problem Statement
Let $\mathcal{M}_{\text{s}}$ be a source LLM with a vocabulary $\mathcal{V}_{\text{s}}$. Given an additional target vocabulary $\mathcal{V}_{\text{new}}$, the aim is to learn a model $\mathcal{M}_{\text{t}}$ with the identical architecture as $\mathcal{M}_{\text{s}}$ that supports a target vocabulary $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$. $\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{t}}$ have their corresponding tokenizer, $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$.
First, a target language-specific auxiliary tokenizer $\mathcal{T}_{\text{aux}}$ is trained on target language data $\mathcal{D}$. $\mathcal{T}_{\text{t}}$ is then derived from $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{aux}}$ along with their corresponding vocabularies $\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{aux}}$ by incorporating top $k$ frequent tokens ($\mathcal{V}_{\text{new}}$) that are in $\mathcal{V}_{\text{aux}}\backslash(\mathcal{V}_{\text{s}}\cap\mathcal{V}_{\text{aux}})$ into $\mathcal{V}_{\text{s}}$, resulting in $\mathcal{V}_{\text{t}}$. Subsequently, $\mathcal{M}_{\text{t}}$ is initialized with the weights of $\mathcal{M}_{\text{s}}$. Its embedding and output layer matrices are expanded such that the former is of dimensionality $|\mathcal{V}_{\text{t}}|\times H_{\text{t}}$ and the latter $H_{\text{t}}\times|\mathcal{V}_{\text{t}}|$, where $H_{\text{t}}$ is the hidden dimensionality of $\mathcal{M}_{\text{t}}$. Each new token representation is initialized by applying a target vocabulary initialization method (§4). If the weights of both matrices are not tied, they are initialized separately. Finally, $\mathcal{M}_{\text{t}}$ is adapted to the target language by training it on $\mathcal{D}$ using a clm objective.
4 Target Vocabulary Initialization
After vocabulary expansion, the new embeddings should be initialized. Our aim is to investigate the effectiveness and robustness of different initialization approaches in low-resource settings. For that purpose, we evaluate: (1) random initialization; (2) initialization based on auxiliary models; and (3) heuristic-based initialization.
4.1 Random Initialization
We randomly initialize new tokens by sampling from $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma$ are the mean and standard deviation of token embeddings of $\mathcal{M}_{\text{s}}$ (Random). This is the most simple and common approach when adapting English-centric LLMs to a target language via vocabulary expansion in high-resource settings Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024).
4.2 Initialization Based on Auxiliary Models
FOCUS (Dobler and de Melo, 2023) is a state-of-the-art CVA method that relies on auxiliary embeddings, i.e. fastText Bojanowski et al. (2017), for initialization. [^3] Due to resource constraints, we only report results using FOCUS as a representative initialization method that relies on auxiliary embeddings since it outperforms other similar methods such as WECHSEL Minixhofer et al. (2022) for vocabulary replacement. The main assumption is that semantic transfer of embeddings should result in better initialization over Random. We apply FOCUS by tokenizing $\mathcal{D}$ using $\mathcal{T}_{\text{t}}$ and train a fastText model for each language. [^4] Preliminary experiments with FOCUS using off-the-shelf pre-trained word-level fastText models yielded lower performance (see Appendix E.2).
4.3 Heuristic-based Initialization
Sophisticated methods such as FOCUS require auxiliary embeddings trained in the target language, which might not be available or hard to train in low-resource settings. Motivated by this, we evaluate a set of heuristic-based initialization methods that do not rely on any external data or model and can be applied to any language.
Mean.
A straightforward approach is to initialize each new token in $\mathcal{V}_{\text{new}}$ with the average of its corresponding source tokens obtained by using $\mathcal{T}_{\text{s}}$ Yao et al. (2021). Koto et al. (2021) and Gee et al. (2022) have also followed a similar approach for vocabulary replacement in domain adaptation.
Merge.
Mean uses solely $\mathcal{T}_{\text{s}}$, which might produce subtokens from $\mathcal{V}_{\text{s}}$ that are not semantically related to the new target token. To overcome this issue, we propose using merge rules from $\mathcal{T}_{\text{t}}$ to effectively initialize $\mathcal{V}_{\text{new}}$. Merge rules describe how $\mathcal{T}_{\text{t}}$ can combine two subtokens into one, e.g. (‘a’, ‘t’) $\rightarrow$ ‘at’. Therefore, each new token in $\mathcal{V}_{\text{new}}$ can be decomposed into several existing subtokens from $\mathcal{V}_{\text{s}}$. We hypothesize that such subtokens can be more semantically related to the new token than those in Mean as they use information from $\mathcal{T}_{\text{t}}$.
We first identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$. Subsequently, we generate a mapping from each new target token to source subtokens using the identified merge rules. Finally, for each new token, we compute the hierarchical mean of the associated source subtoken embeddings using the mapping information. Algorithm 1 in Appendix §A describes our method.
Align.
The tokenization of a given word can change dynamically depending on the context, e.g. ‘the’ can be tokenized as (‘th’, ’e’) or (‘t’, ‘he’). Likewise, a new token in $\mathcal{V}_{\text{t}}$ is not always tokenized in the same way by $\mathcal{T}_{\text{s}}$ when operating on a sequence rather than a single token, which is the case for Mean. To obtain a more fine-grained semantic representation of a new token, we propose token alignment initialization that leverages mapping information between tokens tokenized with $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$ and the mapping frequency (i.e. more frequent mappings are more important for the final representation). This process allows us to consider different tokenization variants for the initialization of a given token.
For each sentence $x\in\mathcal{D}$, we first tokenize it using $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$. We then compare the two tokenized versions and generate a mapping list of tuples for each new token in $x$, e.g. a token $t_{5}$ in $\mathcal{V}_{\text{t}}$ is tokenized as [($t_{1}$, $t_{2}$), ($t_{3}$, $t_{4}$)] in $\mathcal{V}_{\text{s}}$. Subsequently, we concatenate all mapping lists of tuples in $\mathcal{D}$ for each new token $t\in\mathcal{V}_{\text{new}}$. For each $t$, we then construct a unique set of its constituent tuples, i.e. $\bm{T}_{t}=(\bm{t}_{1},\bm{t}_{2},\bm{t}_{3},\dots)$, and the corresponding frequency list $\bm{F}_{t}=(f_{\bm{t}_{1}},f_{\bm{t}_{2}},f_{\bm{t}_{3}},\dots)$, given the concatenated list in the previous step. Finally, we use $\bm{T}_{t}$ and $\bm{F}_{t}$ to aggregate the mapping information for token $t$ and generate its representation by computing $\sum_{\bm{t}\in\bm{T}_{t}}{\left[f_{\bm{t}}\frac{1}{|\bm{t}|}\sum_{t^{\prime}\in\bm{t}}{\bm{e}_{t^{\prime}}}\right]}$, where $\bm{e}_{t^{\prime}}$ is the vector representation (embedding or LM head) of the subtoken $t^{\prime}\in\mathcal{V}_{s}$.
5 Training Strategy
Continual pre-training on $\mathcal{D}$ is an integral part of vocabulary expansion to improve the alignment of the newly initialized embeddings. To this end, we explore: (i) the training procedure, (ii) the objective function, and (iii) the input sequence length.
Procedure.
We use by default low-rank adaptation (LoRAHu et al. (2022) applied to all linear layers, following previous work on vocabulary expansion in high-resource settings Cui et al. (2023); Balachandran (2023); Abbasi et al. (2023); Choi et al. (2024).
We also consider the two-stage tuning process (2-stage) Cui et al. (2023), where only the embeddings and LM head are first updated, followed by tuning the LoRA modules. This process can help minimize the risk of overfitting to the initial embedding state, which might be suboptimal Downey et al. (2023). Alternatively, we train only the top and bottom two layers (2x2 LS) following Remy et al. (2024). This approach aims to calibrate only the parts closely related to the encoding and decoding of the target language, minimizing changes to the base model. Note that none of these approaches have ever been compared to the standard LoRA or against each other. We tune the embeddings and LM head in each case.
Objective.
We first evaluate clm, which has been used by previous work in high-resource settings (Cui et al., 2023; Balachandran, 2023; Larcher et al., 2023, inter alia), as a default training objective. We also consider a multi-token prediction (mtp) objective Gloeckle et al. (2024), where the model must predict multiple consecutive tokens at each timestep rather than a single token. mtp has been proposed for pre-training and exhibited performance gains over clm. However, it has yet to be explored for continual pre-training for cross-lingual transfer. We experiment with one additional LM head (i.e. predicting two consecutive tokens at a time) and initialize the additional weights with those from the original LM head. [^5] We find in our preliminary analysis that random initialization does not work well (Appendix E.3).
Sequence Length.
Finally, we shorten the default input sequence length from 2,048 to 512, thereby increasing the number of training batches. We hypothesize that this is critical in low-resource settings, as a small number of model updates could be prone to underfitting.
6 Experimental Setup
6.1 Source Models
We use Llama2 (7B) Touvron et al. (2023), an English-centric model, with its $\mathcal{T}_{\text{s}}$ based on byte-fallback Byte Pair Encoding (BPE) Sennrich et al. (2016) and $|\mathcal{V}_{\text{s}}|$ set to 32K in the experiments. We also use Llama3 (8B) Dubey et al. (2024) and Gemma2 (9B) Riviere et al. (2024) for consistency in our analysis. These models have a far larger vocabulary size than Llama2, i.e. 128K and 256K, respectively.
6.2 Target Languages and Data
We experiment with a typologically diverse set of ten languages with various scripts. This includes German (Indo-European) and Swahili (Niger–Congo) for the Latin script, and Arabic (Afroasiatic), Burmese (Sino-Tibetan), Greek (Indo-European), Hindi (Indo-European), Japanese (Japonic), Sinhala (Indo-European), Telugu (Dravidian), and Thai (Kra–Dai) for non-Latin scripts. We select these languages because of the availability of downstream task datasets with the same task formulation across languages.
To simulate a realistic low-resource adaptation scenario, we follow Yong et al. (2023) in using 30K sentences per language. These sentences are randomly sampled from their language-specific subcorpus of CC-100 Conneau et al. (2020).
Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
Afroasiatic Sino-Tibetan Indo-European Indo-European Indo-European Japonic Indo-European Niger–Congo Dravidian Kra–Dai
Model mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum
Source .08.03 360.0 .03.00 240.1 .25.01 250.1 .08.03 250.2 .12.03 390.2 .03.03 220.1 .05.00 260.1 .01.00 270.4 .06.00 210.2 .06.02 220.2
LAPT .18.00 360.2 .11.00 240.3 .20.00 210.1 .21.00 250.3 .21.00 380.2 .08.00 230.2 .07.00 274.1 .12.00 300.1 .09.00 290.1 .12.00 230.2
+Speedup Random .07.00 320.3 .05.00 170.3 .28.00 230.3 .13.00 200.1 .09.00 350.2 .18.00 220.2 .07.00 290.1 .13.00 290.1 .06.00 260.2 .07.00 190.2
FOCUS .07.00 320.2 .03.00 160.3 .28.00 220.4 .17.00 180.1 .10.00 340.2 .19.00 220.2 .05.00 290.3 .18.00 290.1 .05.00 260.2 .06.00 170.1
Mean .06.00 330.1 .04.00 200.4 .24.01 230.2 .14.00 190.1 .12.00 340.2 .20.00 230.1 .06.00 300.3 .12.00 290.1 .06.00 270.1 .09.00 220.1
Merge .07.00 330.1 .04.00 90.5 .25.00 230.2 .15.00 180.3 .10.00 340.2 .18.00 230.3 .09.00 300.2 .13.00 290.1 .05.00 260.3 .07.00 210.2
Align .06.00 330.1 .04.00 150.7 .26.01 210.2 .16.00 170.2 .13.00 350.2 .20.00 230.1 .08.00 300.3 .17.00 290.0 .06.00 270.2 .07.00 220.2
Table 1: Mean performance on mt and sum tasks over five runs in low-resource settings (30K sentences) using Llama2 as source. The baselines without vocabulary expansion are in gray. Darker blue and red shades indicate higher positive and negative relative performance change over Source per language and task, respectively.
6.3 Baselines
We use the following two methods as our baselines: (1) the source model $\mathcal{M}_{\text{s}}$ off-the-shelf without any adaptation (Source); (2) we continue pre-training Source on $\mathcal{D}$ by keeping $\mathcal{V}_{\text{s}}$ (LAPT). These do not offer any inference speedups.
6.4 Tasks
We evaluate all models using two tasks per target language in a zero-shot setting: (1) English-to-target machine translation (mt) using FLORES-200 NLLB Team et al. (2022) and (2) summarization (sum) including German MLSUM Scialom et al. (2020), GreekSUM Evdaimon et al. (2023), and XL-Sum Hasan et al. (2021) for the rest.666sum performance is not directly comparable between different languages as the data does not match across languages. Following Ahia et al. (2023), we use 500 random samples. The prompt templates can be found in Appendix B.
6.5 Evaluation Metrics
Task Performance.
We use chrF Popović (2015) for mt and ROUGE-L Lin (2004) for sum. We report average performance over five different runs throughout the experiments and analysis.
Perplexity.
We report perplexity on the 100K language-specific held-out sentences sampled from the CC-100 subcorpus as an auxiliary metric for evaluating model performance.
Inference Efficiency.
We measure inference efficiency as the number of tokens generated per second Hong et al. (2024).
6.6 Implementation Details
We set $|\mathcal{V}_{\text{aux}}|$ to 50K across languages and the number of new target tokens $k$ to 100 by default. We investigate the effect of varying $k$ in §7.4. For LAPT, we train each model for two epochs with a batch size of 8, a maximum learning rate of 1e-4, and a sequence length of 2,048. We set a LoRA rank to 8 following previous work Cui et al. (2023); Abbasi et al. (2023); Lin et al. (2024).
7 Results and Analysis
7.1 Target Vocabulary Initialization
We analyze the effect of different target vocabulary initialization methods (§4) on task performance and inference efficiency using Llama2 as source.
Model ar my de el hi ja si sw te th
Source 8.3 5.0 35.7 4.8 6.4 20.4 3.5 47.2 2.4 9.4
LAPT 4.2 2.7 10.9 2.9 3.2 5.4 2.3 12.7 1.8 4.3
Random 11.9 11.9 12.0 7.1 8.3 15.0 8.7 13.9 7.1 8.8
FOCUS 11.5 13.8 12.1 6.6 8.9 14.8 9.3 13.7 7.9 9.4
Mean 9.4 11.6 11.7 6.0 6.4 14.7 8.1 13.5 6.9 7.8
Merge 9.8 12.6 11.8 6.1 6.7 15.1 8.7 13.7 7.3 8.0
Align 9.3 11.2 11.7 5.9 6.3 15.1 8.0 13.5 6.7 7.6
Table 2: Perplexity on language-specific held-out dataset using Llama2 as source. Note that results are not comparable between models with gray and others due to their difference in vocabulary. Bold and underlined indicate the best and second-best perplexities among adapted models for each language.
Model Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum
42.2 20.1 28.6 15.0 42.6 24.8 42.4 13.3 42.2 21.5 42.8 21.9 28.1 14.8 43.0 24.1 27.7 10.0 42.6 17.9
Source 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x
LAPT 1.01x 1.13x 0.98x 1.01x 0.97x 1.05x 1.00x 0.97x 1.00x 1.01x 0.98x 1.03x 1.00x 1.00x 0.98x 1.21x 0.98x 0.99x 0.98x 1.01x
Random 1.17x 1.88x 1.18x 1.48x 0.95x 0.97x 1.32x 1.79x 1.19x 1.57x 1.04x 1.13x 1.56x 2.54x 0.92x 1.27x 1.93x 4.26x 1.06x 1.55x
FOCUS 1.14x 1.79x 1.13x 1.38x 0.89x 0.94x 1.41x 1.66x 1.11x 1.49x 1.04x 1.11x 1.13x 2.40x 0.93x 1.26x 1.79x 4.26x 1.06x 1.44x
Mean 1.08x 1.99x 1.02x 2.15x 0.95x 0.95x 1.32x 1.73x 1.15x 1.70x 1.08x 1.13x 1.29x 2.67x 0.91x 1.25x 1.84x 4.58x 1.10x 1.62x
Merge 1.17x 1.95x 1.08x 1.20x 0.94x 1.00x 1.40x 1.70x 1.07x 1.57x 1.05x 1.12x 1.58x 2.49x 0.92x 1.25x 1.71x 4.38x 1.05x 1.59x
Align 1.10x 1.98x 1.08x 1.90x 0.95x 1.04x 1.37x 1.67x 1.15x 1.77x 1.09x 1.13x 1.47x 2.58x 0.93x 1.26x 1.84x 4.42x 1.08x 1.63x
Table 3: Inference efficiency measured by the number of tokens generated per second (T/S). The first row of Source corresponds to its respective T/S for each language and task. Bold and underlined indicate the best and second-best speedups across tasks and models with the same base model for each language.
Task Performance and Perplexity.
Table 1 shows the performance of all methods, while Table 2 shows the corresponding perplexities on the held-out language-specific dataset. We first find that models adapted with Mean and Align show comparable or better performance than Source in the majority of the cases. Specifically, they outperform or perform the same as Source in 13 and 14 out of 20 cases, respectively. This result aligns with the perplexity results in Table 2, where Align tends to exhibit the lowest perplexity across languages, followed by Mean. Further, some performance gaps in the rest of the cases are within only a few points ($\leq$ 3). In contrast, Merge shows worse performance than Align and Mean, outperforming or performing the same as Source in 11 cases. We speculate that Merge might be less informative than these two methods, as it does not rely on surface information for initialization, resulting in lower performance.
Random performs the same as or better than Source in 12 cases, while FOCUS achieves this in 11 cases—the worst of the adapted models along with Merge. This is also evident in the perplexity results, where FOCUS shows the worst perplexity in six languages (i.e. my, de, hi, si, te, and th). These results suggest that sophisticated initialization approaches do not always work well, and the popular Random is not the best choice in low-resource settings. Thus, we conclude that models adapted with Mean and Align are more likely to perform competitively to Source.
Next, we analyze the results by language and task. We see that the adapted models, especially Mean and Align, generally perform competitively with Source within a few points of degradation across all languages and substantially outperform it in Greek, Japanese and Swahili mt, with up to 17 points of improvement in Japanese. However, they often do not perform competitively in Burmese, German, Greek, and Hindi sum. We hypothesize that sum might require more training data than mt to perform well as generating long text (up to 128 generated tokens and using long context) requires models to have strong generative capabilities in a target language. We later address this challenge in §7.2 by showing that these performance gaps can be drastically narrowed using our alternative training strategies, further contributing to the competitiveness of vocabulary expansion approaches to Source.
Inference Efficiency.
Table 3 shows the inference efficiency of adapted models across tasks and languages. We first see that FOCUS tends to exhibit the worst speedups across languages in sum, followed by Merge. Specifically, its speedups are the worst in 8 out of 10 languages. These results are consistent with downstream performance and may be due to an inability to effectively generate newly added target tokens, which is essential for improving inference efficiency in a target language. For mt, we generally see smaller differences between models since its input is mostly in English, and thus, only the output contributes to inference speedups.
Looking at the results by language and script, we observe that non-Latin script languages, except for Japanese, show substantial speedups of up to 4.58x. Japanese, on the other hand, exhibits moderate inference speedups of up to 1.13x. This is possibly due to the inclusion of Japanese (0.1%) and Chinese [^7] Japanese shares some scripts (i.e. kanji) with Chinese. data (0.13%) in the Llama2 training corpus Touvron et al. (2023). For a similar reason, German, which accounts for the largest size among non-English languages in the corpus (0.17%), shows no speedups across tasks. Although Swahili is not included in the training corpus, it does not seem to benefit well from vocabulary expansion with Llama2, possibly due to its Latin script (91% of the pre-training data).
Recommendation.
Using Mean or Align can provide better downstream performance and inference speedups. Vocabulary expansion contributes well to inference speedups when a target language is not included in pre-training data and is written in non-Latin scripts. [^8] If inference speedups do not matter, we recommend using LAPT in general. Since it does not involve parameter updates from scratch, it can lead to more stable task performance in low-resource settings.
7.2 Training Strategy
We analyze the effectiveness of each training strategy introduced in §5. Due to limited compute, we only use the best-performing Align as the vocabulary initialization method and experiment with six languages with the largest speedups (i.e. ar, my, el, hi, si, and te) in sum. Table 4 lists the performance of the adapted models.
Arabic Burmese Greek Hindi Sinhala Telugu
Model mt sum mt sum mt sum mt sum mt sum mt sum
Source .08.03 360.0 .03.00 240.1 .08.03 250.2 .12.03 390.2 .05.00 260.1 .06.00 210.2
LAPT .18.00 360.2 .11.00 240.2 .21.00 250.3 .21.00 380.2 .07.00 274.0 .09.00 290.1
LoRA clm+2048 .06.00 330.1 .04.00 150.7 .16.00 170.2 .13.00 350.2 .08.00 300.3 .06.00 270.2
mtp+2048 .11.00 320.2 .03.00 140.7 .14.00 200.2 .12.00 340.2 .08.00 290.1 .04.00 250.1
clm+512 .12.00 330.1 .12.00 230.1 .23.00 190.4 .17.00 350.1 .11.00 320.1 .10.00 280.1
mtp+512 .14.00 330.1 .12.00 230.1 .21.00 210.3 .15.00 350.1 .11.00 300.3 .09.00 280.1
2-stage clm+2048 .09.00 330.1 .04.00 170.2 .13.00 210.1 .13.00 360.1 .09.00 300.4 .05.00 280.1
mtp+2048 .13.00 330.1 .02.00 170.3 .19.00 210.1 .11.00 360.1 .08.00 290.2 .03.00 250.2
clm+512 .12.01 330.1 .07.00 130.5 .22.00 220.3 .16.00 350.1 .10.00 310.3 .07.00 280.1
mtp+512 .11.00 330.2 .07.00 170.2 .21.00 220.1 .16.00 350.1 .08.00 310.3 .07.00 280.1
2x2 LS clm+2048 .11.01 330.1 .14.00 250.1 .24.00 220.3 .15.00 360.2 .11.00 320.2 .11.00 290.1
mtp+2048 .15.01 330.1 .11.00 230.2 .23.00 230.2 .12.00 360.0 .12.00 320.1 .12.00 280.2
clm+512 .11.00 330.1 .15.00 260.2 .25.00 230.3 .19.00 360.1 .11.00 330.2 .13.00 300.1
mtp+512 .17.00 330.2 .16.00 260.2 .24.00 230.2 .15.00 360.1 .13.00 330.3 .12.00 290.1
Table 4: Mean performance of Align models (five runs) with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Looking at the results by training procedure, we observe that 2x2 LS drastically improves performance across tasks and languages, achieving up to 10 points improvement from LoRA (Burmese mt and sum with clm+2048). This suggests that the frequently used full LoRA approach in high-resource settings is not the best training approach in low-resource settings. Although 2-stage does not consistently provide gains over LoRA, especially for mt, it does not show substantial performance degradation either. Next, we look at the effectiveness of mtp. Although it does not always improve performance over clm, it boosts Arabic mt performance across models, especially with 2x2 LS. Similar to 2-stage, its performance degradation, if it exists, is generally within a few points ($\leq 3$). Finally, usign a short sequence length (512) works well across models and languages, showing improvements of up to 9 points (LoRA+mtp in Burmese and 2-stage+clm in Greek mt), confirming our hypothesis (§5) that increasing the number of model updates can help avoid underfitting.
Recommendation.
2x2 LS generally leads to performance improvements. The short training sequence length of 512 can also contribute to performance improvements in low-resource settings. Finally, mtp sometimes improves performance with no major performance drop in other cases.
7.3 Experiments with Other Source LLMs
We investigate whether models other than Llama2 adapted based on our recommendations in §7.1 and §7.2 benefit from inference speedups while maintaining competitive performance to their base models. We use Llama3 and Gemma2 as source. We apply the 2x2 LS+mtp+512 training strategy and initialize models with either Mean or Align along with Random as a baseline. Due to resource constraints, we only experiment with Burmese, Sinhala, and Telugu in the remainder of the paper, as they are the worst fragmented languages of our target languages (see Appendix E.1). Table 5 shows the performance of adapted models using Llama3 and Gemma2 as source.
Burmese Sinhala Telugu
Llama3 mt sum mt sum mt sum
Source .09.00 280.0 .19.00 300.0 .24.00 290.0
LAPT .17.00 280.1 .21.00 310.1 .30.00 290.1
+Speedup Random .19.00 230.2 .14.00 310.2 .13.00 290.1
Mean .19.00 230.2 .12.00 320.1 .12.00 290.2
Align .21.00 280.3 .22.00 330.1 .31.00 310.1
Speedup 2.60x 3.52x 2.36x 3.19x 2.15x 3.55x
Gemma2
Source .21.01 340.1 .19.00 340.2 .31.01 300.2
LAPT .28.00 280.3 .29.00 340.1 .43.00 290.1
+Speedup Random .24.00 260.3 .26.00 330.1 .35.00 280.2
Mean .25.00 250.3 .27.00 320.2 .33.00 280.1
Align .25.00 250.2 .28.00 320.1 .32.00 290.1
Speedup 2.24x 1.53x 2.26x 1.53x 1.03x 1.25x
Table 5: Mean performance and inference speedup over five runs with Llama3 and Gemma2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Overall, models adapted with Align generally perform competitively with Gemma2 Source in Sinhala and Telugu sum tasks and often outperform Llama3 Source in both tasks and Gemma2 Source in mt. Further, Align models exhibit substantial inference speedups of up to 3.55x and 2.26x for Llama3 and Gemma2, respectively. These results confirm the versatility of our strategies regardless of the type of base model. Although models adapted with Mean rival Source, they, along with Random, underperform it in Burmese sum and Sinhala and Telugu mt tasks when using Llama3 as source. This can be due to underfitting, i.e. their perplexities are higher than Align (Appendix E.4).
Examining the results by source model and task, we observe that the adapted models with Llama3 and Gemma2 as source show substantially better mt performance than those with Llama2 despite their similar model size. In particular, the Telugu-adapted models initialized with Align with Llama3 and Gemma2 as source obtain 19 and 20 points better performance than those Llama2 counterparts. This trend suggests a successful cross-lingual transfer of the generative capabilities of the base models. However, sum does not seem to follow the trend, i.e. the adapted models with Llama3 and Gemma2 do not outperform those with Llama2. In fact, their semantic-level performances measured by BLEURT Sellam et al. (2020) improve up to 7 points from those with Llama2 (Appendix E.4). Thus, adapted models actually enjoy their base model capabilities to improve performance at the semantic level rather than the surface level measured by ROUGE-L.
Figure 2: Downstream performance and inference speedup in (a) mt and (b) sum across different $|\mathcal{V}_{\text{new}}|$. Red and gray dotted lines denote Source and LAPT.
7.4 Target Vocabulary Size
The extent to which a target language benefits from inference speedups varies across languages, models, and tasks, as observed from Tables 3 and 5. Although larger $|\mathcal{V}_{\text{new}}|$ can lead to faster inference, it comes with the risk of underfitting in low-resource settings. [^9] Given a small and limited $\mathcal{D}$, the larger $|\mathcal{V}_{\text{new}}|$, the less fragmentation with respect to $\mathcal{D}$, the fewer the number of training tokens, and therefore the fewer model updates. To help select an optimal $|\mathcal{V}_{\text{new}}|$ for obtaining substantial speedups while retaining competitive performance to Source, we conduct a cost-benefit analysis with respect to $|\mathcal{V}_{\text{new}}|$. We experiment with $|\mathcal{V}_{\text{new}}|=\{50,100,500,1\text{K},5\text{K}\}$ while previous studies set $|\mathcal{V}_{\text{new}}|$ to over 10K, e.g. 17,953 for Chinese LLaMA. For comprehensiveness, we employ all three source models and follow the same setup as in §7.3. Figure 2 shows the performance changes and corresponding inference speedups with respect to different $|\mathcal{V}_{\text{new}}|$.
We first observe that the larger $|\mathcal{V}_{\text{new}}|$, the worse the task performance across different source models, languages, and tasks in general. Notably, sum appears to be less robust to the changes in $|\mathcal{V}_{\text{new}}|$ than mt across different models and languages. Adapted models underperform Source even at $|\mathcal{V}_{\text{new}}|=500$ in the majority of the cases, while they, especially when initialized with Align, still outperform or rival Source in mt in almost all the cases. This difference can be due to the difficulty of each task as discussed in §7.1, i.e. sum requires more $\mathcal{D}$ than mt to perform well. These suggest that $|\mathcal{V}_{\text{new}}|$ should be set smaller (e.g. $|\mathcal{V}_{\text{new}}|=100$ in our setup) for tasks like sum to be competitive with Source in task performance.
[…]
Recommendation.
Setting $|\mathcal{V}_{\text{new}}|$ between 500 and 1K for mt and 100 for sum can be a suitable starting point to maintain competitive performance to Source while benefiting greatly from inference speedups under low-resource settings.
8 Conclusion
We investigated cross-lingual vocabulary expansion in low-resource settings across vocabulary initialization approaches and training strategies. Our extensive experiments reveal that a widely used approach in high-resource settings is not always optimal in low-resource settings. In contrast, models adapted by our alternative strategies achieve faster inference while rivaling their base models in task performance. We supplement our analysis with specific recommendations for effective vocabulary expansion. In future work, we will compare vocabulary expansion and vocabulary replacement for efficient LM inference in low-resource settings.
Limitations
Tokenizer.
Heuristic-based target vocabulary initialization with Merge assumes the use of a BPE-based tokenizer which is a common choice in recent LLMs, e.g. Gemma2, Llama3, Llama2, Mistral Jiang et al. (2023), inter alia. Experimenting with other tokenizers, such as Unigram Kudo (2018), is beyond the scope of this paper.
Languages.
Our work covers ten linguistically diverse languages, i.e. Arabic, Burmese, German, Greek, Hindi, Japanese, Sinhala, Swahili, Telugu, and Thai, following previous work on CVA that has also tested a similar number of languages. For instance, Downey et al. (2023) tested ten languages, and Minixhofer et al. (2022) tested eight languages. Experimenting with more languages is an interesting avenue for future work but out of the scope of this paper given our limited computing capacity.
Model Size.
Our experiments consider LLMs of 7B, 8B, and 9B parameters. Investigating the performance of vocabulary expansion with larger models would be valuable in future studies. Limited access to computational resources did not allow us to experiment with larger LLMs.
Tasks.
Our experiments consider two generation tasks, i.e. machine translation and summarization, following Hong et al. (2024). Nonetheless, investigating different types of downstream tasks (e.g. reasoning) would be an interesting direction but out of the scope of this paper, given our limited computation and limited language resources available.
Acknowledgments
We would like to thank Miles Williams, Huiyin Xue, and Constantinos Karouzos for the comments. We acknowledge IT Services at the University of Sheffield for the provision of services for high-performance computing.
Appendix
Appendix A Pseudo-code for Merge Rule-based Initialization
Algorithm 1 describes the pseudo-code of vocabulary initialization using merge rules.
Data: $\mathcal{V}_{\text{s}}$, $\mathcal{V}_{\text{new}}$, Merge rules of $\mathcal{T}_{\text{t}}$: $Merges$, Embeddings of $\mathcal{M}_{\text{t}}$: $\bm{E}$
Result: Updated $\bm{E}$
1
/* Identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$/
2 $Merges_{\text{new}}\leftarrow\text{List}[~{}]$;
3 for $m\in Merges$ do
4       $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
5       $t\leftarrow t_{1}+t_{2}$;
6       if $t\in\mathcal{V}_{\text{new}}$ then
7             $AddToMerges_{\text{new}}(m)$;
8
9       end if
10
11 end for
12
/* Generate a mapping from a new token to source subtokens using $Merges_{\text{new}}$/
13 $Maps\leftarrow\text{Dict}[t\in\mathcal{V}_{\text{new}},\text{List}[~{}]]$;
14 while $|Merges_{\text{new}}|\neq 0$ do
15       $UnmapMerges\leftarrow\text{List}[~{}]$;
16       for $m\in Merges_{\text{new}}$ do
17             $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
18             $t\leftarrow t_{1}+t_{2}$;
19             if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
20                   $AddToMaps[t]([t_{1},t_{2}])$;
21
22            else if $t_{1}\in\mathcal{V}_{\text{new}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
23                   if $Maps[t_{1}]\neq\text{List}[~{}]$ then
24                         $AddToMaps[t]([Maps[t_{1}],t_{2}])$
25                  else
26                         $AddToUnmapMerges(m)$
27                   end if
28
29            else if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{new}}$ then
30                   if $Maps[t_{2}]\neq\text{List}[~{}]$ then
31                         $AddToMaps[t]([t_{1},Maps[t_{2}]])$
32                  else
33                         $AddToUnmapMerges(m)$
34                   end if
35
36             else
37                   if $Maps[t_{1}]\neq\text{List}[~{}]\cap Maps[t_{2}]\neq\text{List}[~{}]$ then
38                         $AddToMaps[t]([Maps[t_{1}],Maps[t_{2}]])$
39                  else
40                         $AddToUnmapMerges(m)$
41                   end if
42
43             end if
44
45       end for
46
47      if $|UnmapMerges|=0$ then
48             break;
49
50       end if
51      $Merges_{\text{new}}\leftarrow UnmapMerges$;
52
53 end while
54
/* Obtain vector representation for each new token using $Maps$/
55 for $t\in\mathcal{V}_{\text{new}}$ do
56       $\bm{E}[t]\leftarrow ComputeHierarchicalMean(\bm{E}[Maps[t]])$;
57 end for
58
Algorithm 1 Merge Rule-based Initialization
Appendix B Prompt Templates
We translate the English prompt templates provided by Ahia et al. (2023) for sum with a machine translation API, following Yong et al. (2023). For mt, we formulate a task-specific English prompt, followed by machine translation for each language. Table 8 shows the prompt templates used in our evaluation.
Table 8: Prompt template for each task and language.
Appendix C Implementation Details
C.1 Language Adaptive Pre-training
Table 6 lists the hyperparameters in LAPT.
C.2 Evaluation
sum.
To compute ROUGE-L, we split sentences with an mT5 Xue et al. (2021) tokenizer as preprocessing following Maynez et al. (2023) and subsequently call rouge_scorer [^10] https://github.com/google-research/google-research/tree/master/rouge to compute the metric. To compute BLEURT, we use BLEURT-20, which is available from https://github.com/google-research/bleurt.
Hyperparameters.
Table 7 lists the parameters used during evaluation. To make a fair comparison, we do not conduct any generation parameter tuning and use the same ones across all approaches. For sum, we truncate an article whenever it exceeds the maximum prompt length of 4,096 to avoid the CUDA out-of-memory error.
C.3 Libraries and Hardware
For Llama2 and Gemma2, we train tokenizers using SentencePiece Kudo and Richardson (2018) and convert them into the Hugging Face Tokenizers Moi and Patry (2023) format. For Llama3, we train tokenizers using Hugging Face Tokenizers. We implement our models using PyTorch Paszke et al. (2019), Hugging Face Transformers Wolf et al. (2020) and PEFT Mangrulkar et al. (2022). We preprocess datasets with Hugging Face Datasets Lhoest et al. (2021). For evaluation, we use Hugging Face LightEval [^11] https://github.com/huggingface/lighteval. We use a single NVIDIA A100 (80GB) for all experiments.
C.4 Code and Models
Our code and models are available on https://github.com/gucci-j/lowres-cve.
Appendix D Licenses
This study uses various publicly available models and datasets with different licenses, as detailed below, all of which permit their use for academic research.
D.1 Models
Llama2 is licensed under the Llama 2 Community License Agreement. [^12] https://ai.meta.com/llama/license/ Likewise, Llama3 is licensed under the Llama 3 Community License Agreement. [^13] https://llama.meta.com/llama3/license/ Gemma2 is licensed under the Gemma Terms of Use Agreement. [^14] https://ai.google.dev/gemma/terms
Figure 3: Average number of tokens on the FLORES-200 dev set across languages and models.
D.2 Datasets
FLORES-200 is distributed under CC BY-SA 4.0. XL-Sum is licensed under CC BY-NC-SA 4.0, while MLSUM is distributed under an MIT license. GreekSUM is not made public, but the official script for reproduction is available on https://github.com/iakovosevdaimon/GreekSUM.
Appendix E Results and Analysis
E.1 Overfragmentation
Figure 3 shows the average number of tokens on the FLORES-200 dev set across languages and models. Following the assumption by Ahia et al. (2023) that “language A is more efficiently tokenized than language B if it uses fewer tokens per sentence on average,” [^15] Note that this assumption is only applicable to the FLORES-200 dataset as it is a parallel machine translation corpus. we can clearly see that (1) Burmese, Sinhala, and Telugu are the worst overfragmented languages among the ten target languages across all models, and (2) English is the most efficiently tokenized language.
E.2 Target Vocabulary Initialization
Table 9 compares models adapted with FOCUS using Llama2 as source in Burmese, Sinhala, and Telugu. Subword-fastText is the exact same model as FOCUS in Table 1 and makes use of only $\mathcal{D}$ to train the fastText model. Word-fastText uses a pre-trained language-specific word-level fastText model provided by Grave et al. (2018), i.e. with access to an external model. We observe that models using a word-level pre-trained fastText model consistently underperform those using a subword-level fastText model except for sum in Sinhala. Therefore, we conclude that using a pre-trained word-level fastText model does not help improve performance for models adapted with FOCUS in low-resource settings.
Burmese Sinhala Telugu
Model mt sum mt sum mt sum
Source .03.00 240.1 .05.00 260.1 .06.00 210.2
LAPT .11.00 240.3 .07.00 274.1 .09.00 290.1
FOCUS
Subword-fastText .03.00 160.3 .05.00 290.3 .05.00 260.2
Word-fastText .01.00 50.2 .04.00 290.3 .03.00 240.2
Table 9: Mean performance of models initialized with different variants of FOCUS over five runs with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Arabic Greek Hindi
Model mt sum mt sum mt sum
Source .08.03 360.0 .08.03 250.2 .12.03 390.2
LAPT .18.00 360.2 .21.00 250.3 .21.00 380.2
LoRA mtp + 2048 .11.00 320.2 .14.00 200.2 .12.00 340.2
mtp $\ominus$ Copy + 2048 .10.00 320.1 .15.00 200.1 .11.00 350.2
2-stage mtp + 2048 .13.00 330.1 .19.00 210.1 .11.00 360.1
mtp $\ominus$ Copy + 2048 .10.01 330.1 .17.00 220.2 .10.00 360.1
2x2 LS mtp + 2048 .15.01 330.1 .23.00 230.2 .12.00 360.0
mtp $\ominus$ Copy + 2048 .09.00 330.1 .15.00 210.1 .09.00 350.1
Table 10: Mean performance of models initialized with Align over five runs with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Model my si te
Source 3.7 3.5 3.0
LAPT 1.9 2.0 1.8
Random 6.9 7.1 6.1
Mean 6.7 6.8 5.8
Align 6.0 5.5 4.6
(a) Llama3
Model my si te
Source 48.0 59.9 51.9
LAPT 4.6 4.3 5.2
Random 9.0 6.9 6.7
Mean 9.0 6.9 6.6
Align 9.0 6.9 6.6
(b) Gemma2
Table 11: Perplexity on language-specific held-out dataset using Llama3 and Gemma2 as source. Note that results are not comparable between models with gray and others due to their difference in vocabulary.
E.3 Training Strategies
Table 10 briefly compares mtp approaches with and without additional LM head initialization based on the original LM head weights in Arabic, Greek, and Hindi. $\ominus$ indicates that an additional LM head is randomly initialized. We can see that although models without additional LM head initialization based on their original LM head weights show similar performance to those counterparts in sum, they generally do not outperform those in mt. Therefore, we conclude that it is essential to initialize the weights of an additional LM head with those from the original LM head for effective continual pre-training under low-resource settings.
Burmese Sinhala Telugu
Llama2 ROUGE-L BLEURT ROUGE-L BLEURT ROUGE-L BLEURT
Source 240.1 .03.00 260.1 .19.00 210.2 .13.00
LAPT 240.3 .04.00 27.4.1 .25.00 290.1 .22.00
+Speedup Random 240.1 .07.00 320.2 .26.00 290.1 .23.00
Mean 250.1 .06.00 320.1 .25.00 290.1 .23.01
Align 260.2 .08.01 330.3 .26.01 290.1 .24.00
Llama3
Source 28.00 .06.00 30.00 .24.00 29.00 .22.00
LAPT 280.1 .06.00 310.1 .25.00 300.1 .29.00
+Speedup Random 230.2 .05.00 310.2 .27.00 290.1 .24.01
Mean 230.2 .05.00 320.1 .29.01 290.2 .26.00
Align 280.3 .07.00 330.1 .31.00 310.1 .28.00
Gemma2
Source 340.1 .10.00 340.2 .32.00 300.2 .30.00
LAPT 280.3 .10.00 340.1 .32.01 290.1 .26.00
+Speedup Random 260.3 .10.00 330.1 .34.01 280.2 .26.00
Mean 250.3 .09.00 320.2 .33.00 280.1 .27.00
Align 250.2 .10.00 320.1 .33.01 290.1 .28.00
Table 12: Mean sum performance in ROUGE-L and BLEURT over five runs. Darker blue and red indicate higher positive and negative relative performance change over Source.
E.4 Experiments with Other Source LLMs
Perplexity.
Table 11 shows the perplexities measured on the language-specific held-out dataset using Llama3 and Gemma2 as source.
ROUGE-L vs. BLEURT.
Table 12 shows the sum performances measured by ROUGE-L and BLEURT across models when using the 2x2 LS + mtp + 512 strategy and the Align initialization. As discussed in §7.3, while performances at the surface level measured by ROUGE-L do not substantially change across different base models, those at the semantic level measured by BLEURT change, especially between Llama2 and Gemma2.
E.5 Target Vocabulary Size
Figure 4 shows the average target token ratio across models, tasks, and languages with respect to different $|\mathcal{V}_{\text{new}}|$.
Llama2.
[…]
Llama3.
We observe that the ratio of target tokens for outputs plateaus at around $|\mathcal{V}_{\text{new}}|=500$ in both tasks. After that, there are some drops in the target token ratio in mt. Interestingly, the ratio of target tokens for outputs is quite similar to that for inputs in sum. suggesting that adapted models effectively utilize newly added tokens in outputs. To further optimize inference efficiency, we hypothesize that adding target tokens with a larger number of characters might work. We leave this verification for future work.
Gemma2.
[…]
Moreover, the ratio of target tokens is far lower than that of Llama2 and Llama3, indicating that newly added tokens are underutilized in Gemma2. This could possibly be due to byte-fallback BPE used by Gemma2. Considering Llama3 (i) employs byte-level BPE, (ii) shows substantial speedups with vocabulary expansion in Table 5 and Figure 2, (iii) has also a large vocabulary size of 128K, and (iv) has similar fragmentation rates to Gemma2 for the three languages in Figure 3, vocabulary expansion might work better with byte-level BPE rather than byte-fallback BPE. We hypothesize that since byte-level BPE operates on a byte sequence rather than a raw text, it is, therefore, not affected by language scripts. This nature can help to integrate both a new vocabulary and merge rules with the corresponding source ones. We leave a thorough investigation of this hypothesis for future work.
Figure 4: Average target token ratio in input ($\times$) and output ($\bullet$) with respect to $|\mathcal{V}_{\text{new}}|$ across models, languages, and tasks.